### v1

# Visual Riddles: a Commonsense and World Knowledge Challenge for Large Vision and Language Models

Abstract
Imagine observing someone scratching their arm; to understand why, additional context would be necessary. However, spotting a mosquito nearby would immediately offer a likely explanation for the person’s discomfort, thereby alleviating the need for further information. This example illustrates how subtle visual cues can challenge our cognitive skills and demonstrates the complexity of interpreting visual scenarios. To study these skills, we present Visual Riddles, a benchmark aimed to test vision and language models on visual riddles requiring commonsense and world knowledge. The benchmark comprises 400 visual riddles, each featuring a unique image created by a variety of text-to-image models, question, ground-truth answer, textual hint, and attribution. Human evaluation reveals that existing models lag significantly behind human performance, which is at 82% accuracy, with Gemini-Pro-1.5 leading with 40% accuracy. Our benchmark comes with automatic evaluation tasks to make assessment scalable. These findings underscore the potential of Visual Riddles as a valuable resource for enhancing vision and language models’ capabilities in interpreting complex visual scenarios.
Figure 1: Introducing Visual Riddles, designed to test models on their ability to use commonsense, world knowledge, hints, attributions, and factuality in interpreting complex visual cues. This resource aims to enhance models capability to handle nuanced and factual visual scenarios.
1 Introduction
Humans intuitively utilize commonsense reasoning to interpret complex elements in visual scenes, a skill that current vision and language models frequently lack. For instance, the simple act of a person scratching their arm gains added significance when a mosquito is present on a nearby nightstand, as depicted in Fig. 1. While humans easily recognize such contextual nuances, existing image-understanding models struggle to integrate visual cues with world knowledge stemming from cultural aspects, life-experiences, and physical or social knowledge [1, 2, 3, 4]. This gap has spurred the development of various benchmarks like VisIT-Bench [5] and Encyclopedic VQA [6], which rely on pre-existing images to formulate challenging questions. While effective, this approach risks using images seen during pre-training of large vision-language models and restricts the scope of scenarios to those already captured, potentially limiting creativity and challenge variety.
To address these shortcoming, we introduce Visual Riddles, a benchmark comprising 400 visual riddles, each featuring a question and a synthetic image generated specifically for the challenge. The process of creating a riddle involves designing a scenario with embedded clues that appear as natural elements, then translating this scenario into both an image and a question (§2). For instance, in the top example of Fig. 1, a seemingly inconspicuous mosquito acts as a crucial clue for the source of the person’s discomfort. This method, going from scenario to image, and not the other way around, not only enhances creative liberty but also broadens the range of everyday situations explored, challenging both humans and models in their interpretative and commonsense reasoning abilities. The benchmark also incorporates textual hints and attributions (see  Fig. 1) to direct focus to clues and provide external sources of verification for answers, thus expanding the challenge’s scope and depth.
Our benchmark’s main task (§3) involves solving riddles in an open-ended visual question answering (VQA) format, which takes as input an image and a question, and expects a free-text answer. This setup evaluates the ability to detect subtle visual cues and apply commonsense reasoning or world knowledge to formulate answers. Additionally, we investigate the impact of including hints and attributions in the input to enhance comprehension.
Yet, as in every QA benchmark, evaluation is a key challenge in this setting. To facilitate scalable research, we introduce two additional tasks. The first is a multiple-choice version of the main task, also for the hint- and attribution-assisted variants, allowing for easy automatic scoring based on accuracy. The second task assesses the ability of models to determine the accuracy of open-ended responses in two settings: reference-free, where models evaluate responses based only on the image and the question, and reference-based, where the correct answer is also provided. This task suggests auto-raters to evaluate our riddles, and aim to advance research on such automatic evaluation methods.
Experimental results (§4) reveal a significant gap in performance between humans and state-of-the-art vision language models, with the top-performing model, Gemini-Pro-1.5 [7], only achieving 40% accuracy versus humans’ 82%. Surprisingly, the multiple-choice format proved almost as difficult, with performances slightly superior to those in the open-ended task, but with notably high performances when auxiliary data is included. Automated tests, both reference-free and reference-based, show that Gemini-Pro-1.5, the top auto-rater, matches human judgment 87% of the time in reference-based evaluations, propose a suitable method for evaluating open-ended answers. We also find that some models fail to respond to visual clues accurately when tasked with evaluating answers’ validity (§5.1). Finally, attempts to reproduce the benchmark’s images, rich in nuanced visual clues, using the same prompts and various text-to-image models (§5.2), result in a mere 15% success rate, showcasing the unique challenges visual riddles present to current generative models.
Overall, the findings suggest that the Visual Riddles poses a significant challenge even to state-of-the-art vision-and-language models, emphasizing the critical need for further developments in commonsense reasoning and world knowledge integration to improve model performance on complex visual riddles. To support future research and model evaluation, we make the Visual Riddles dataset, code, and leaderboard publicly available at https://visual-riddles.github.io/.
Figure 2: Overview of the Visual Riddles tasks: (1) Main Task: Solve open-ended questions. (2) Utilizing Hints: Use textual aids to identify key visual clues in riddles. (3) Employing Attributions: Apply web-sourced attributions to improve world-knowledge. (4) Multiple Choice: Select the correct answer to the riddle from five options. (5) Automatic Evaluation: Evaluate open-ended answers in two scenarios— Reference-Free, assessing the correctness of a response based only on the visual riddle, and Reference-Based, comparing responses to the ground truth answer.
2 Data Collection
The Visual Riddles Challenge uses visual clues in images to test common-sense reasoning in vision-and-language models. The goal is to develop a dataset with images that represent ambiguous and culturally rich scenarios. To solve the riddles, models must detect subtle visual clues and engage in reasoning that integrates commonsense and world knowledge.
The Visual Riddles benchmark was hand-curated by seven designers experienced in generating images with text-to-image models. The design process was guided by comprehensive instructions to create visual riddles that were complex enough to challenge models yet solvable by humans. Each riddle consisted of a synthetic image paired with a question and a corresponding ground truth answer (see examples in Fig. 1). Further elaboration on the image generation process was in A.2. To generate high-quality images, the designers had access to advanced text-to-image models, including Midjourney, Ideogram, Canva, Gemini [8], SDXL [9], DALL-E 3 [10], and Stable-Diffusion [11]. The generated images included subtle clues crucial for solving the riddles, and designers were encouraged to embed cultural nuances and their own world knowledge into the riddles. The images were not only photo-realistic but also contained all necessary elements to lead to the correct answer. Designers provided additional hints to the visual clues that guided where to look in the image when trying to solve the riddle (see top example in Fig. 1) and, for riddles requiring world knowledge, included attributions to relevant sources (see bottom example in Fig. 1). After creation, each riddle was peer-reviewed by at least three other designers to ensure the hint’s clarity and the riddle’s solvability. Upon approval, the designer drafted a detailed answer that logically explained the solution based on the visual clues. The Visual Riddles benchmark is licensed under Apache License 2.0.
Commonsense and World Knowledge Categorization and Difficulty Analysis
Finally, we categorize the instances based on the type of commonsense reasoning and knowledge required, including safety-related knowledge, biological knowledge, cultural knowledge, and many more (see A.3). Each riddle is assigned to one of 16 distinct categories, labeled with the single category that best fits it. Additionally, each riddle is assigned a Difficulty Level Index, which quantifies its complexity, ranging from simple, straightforward clues to obscure ones requiring specialized knowledge.
3 Visual Riddles Benchmark
This study introduces three vision-and-language tasks within the Visual Riddles benchmark to evaluate model capabilities. The tasks are: solving open-ended visual riddles, choosing the correct answer from multiple options, and conducting automated evaluations on open-ended riddles. These tasks may incorporate auxiliary information such as hints or attributions to enhance assessment accuracy.
Open-ended VQA
In this task, models are evaluated on their ability to solve visual riddles comprising of an image and a question, requiring not only correct solutions but also detailed explanations. Questions may be binary, requiring a ‘yes’ or ‘no’ answer with justification, or open-ended, inviting freely-formulated responses to queries like ‘why’, ‘where’, or ‘how’. These questions require both locating visual clues in the images and integrating these clues with commonsense reasoning and world knowledge to formulate coherent and correct answers. For example, in Fig. 2.1, a good answer to the question “What is probably the gender of the cat?” should reference both the visual clue (the cat’s fur color) and the relevant world knowledge (Calico cats are predominantly female).
Multiple-choice VQA
We propose this task as an alternative to the open-ended VQA, which can be evaluated automatically using a simple accuracy metric. It involves selecting the correct answer from five options for a visual riddle comprising an image and a question, as shown in Fig. 2.4. Details on how these candidate answers were collected are provided in §4.3.
Open-ended VQA Automatic Evaluation
In this task, models are being assessed for their ability to evaluate the accuracy of open-ended responses to visual riddles. As outlined in Fig. 2.5, it has two categories: (a) Reference-Free, where the model assesses a candidate answer based only on the image and question; and (b) Reference-Based, where it also considers the ground-truth answer to evaluate the candidate’s response. This framework aims to identify the best auto-rater for open-ended tasks, supporting automated leaderboard creation. Designed for scalable evaluation of open-ended responses, this task uses vision and language models to judge both Reference-Free and Reference-Based responses. The Reference-Free baseline is included despite available ground truths, as it sometimes outperforms the Reference-Based approach [5], enhancing our insights into model performance across different contexts.
Auxiliary Information
We utilized auxiliary information within tasks of our benchmark to assess its impact on models performances. Specifically, we incorporate textual hints and source attributions to enhance the model’s ability to solve visual riddles by providing targeted contextual cues. Textual hints are concise directives that focus the model’s attention on specific visual clues within the image. For instance, a hint such as “Look at the colors of the fur” in the calico cat example (see Fig. 2.2) effectively directs the model’s focus towards the type of the cat. Additionally, attributions provide essential world knowledge through a webpage URL from which we extract all the text content, offering the models the necessary information to solve the visual riddle. For example, providing a text stating that calico cats are “almost exclusively female” (see Fig. 2.3) assists models in more effectively inferring a cat’s gender from its fur color.
Figure 3: Amazon Mechanical Turk interface for selecting answers to open-ended riddles. Annotators are presented with an image, a question and several candidate answers, including both human-curated and model-generated predictions, and are tasked with identifying the correct responses.
4 Experiments
We evaluate several state-of-the-art publicly accessible vision-and-language models on each of the tasks in Visual Riddles, which we overview in §3. Then, we describe the experimental setup and results of each of the benchmark tasks, including open-ended VQA (§4.2), multiple-choice VQA (§4.3) and open-ended VQA automatic evaluation (§4.4). For riddles evaluated with auxiliary data, which may have lengthy prompts, we only test models capable of accommodating such lengths.
4.1 Models
We evaluate LLaVA-1.5-7B [12], LLaVA-1.6-34B [13], InstructBLIP-7B [14], GPT4-turbo-preview [15], Gemini-Pro-Vision [8], and Gemini-Pro-1.5 [7]. We used the most up-to-date versions of each model, as of May 2024. Each experiment required prompts of varying token lengths; notably, the prompts for attribution tasks were particularly lengthy, as they also included the attributing texts. Models selection is therefore based on their capacity to accommodate the full prompt required for each task, ensuring that inputs are processed fully without truncation. For further details, see A.4.
4.2 Open-ended VQA
We start our investigation by assessing the performance of humans and models on the primary open-ended VQA task, which requires solving riddles based solely on the accompanying images. To that end, we collect human responses to the riddles through crowdsourcing, as well as prompting several vision-language models, and evaluate all responses using human judgement.
Human Answers
We utilized Amazon Mechanical Turk to collect human open-ended answers for the benchmark’s riddles. Rigorous qualification ensured only 10 highly-rated workers, previously experienced in similar tasks, annotated each of the 400 riddles (three workers per riddle). During annotation, we instructed annotators to not only answer the question but also provide justifications. Annotators were also encouraged to utilize tools like Google Lens and different search engines to research clues and ensure accurate identification of objects. For example, in Fig. 2.4, annotators which are unfamiliar with the Kakapo bird were advised to use Google Lens to identify it. For more details on the annotation process, including full guidelines and UI screenshot examples, see A.5.
Model Answers
[…]
Human Rating of Responses
We assess the correctness of human and model answers using three annotators in a multiple-choice selection format. They select responses that match the ground truth, without any hallucinations, based on provided images, questions, and ground-truth answers. The final rating is determined by a majority vote. Annotator agreement, measured by Krippendorff’s alpha [16], reached 79%. An example of human annotation from Visual Riddles is shown in Fig. 3. Further details on annotation guidelines and results can be found in A.6.
Results
[…]
Table 1: Human ratings for large vision and language models, caption generation to large language models pipelines, and human evaluators on 400 riddles in the Visual Riddles benchmark.
% Human Rating $\uparrow$
LVLM Gemini Pro 1.5 40
Gemini Pro Vision 30
GPT4 34
LLaVA-1.6-34B 15
LLaVA-1.5-7B 13
InstructBlip 13
[…]
Human (Oracle) $\rightarrow$ Gemini Pro 1.5 50
Humans 82
4.3 Multiple-choice VQA
We next evaluate the performance of large vision and language models on the multiple-choice VQA task of the Visual Riddles benchmark, shifting from a generative to a classification task. Each riddle includes an image and a question, along with five answer choices: the correct one from the designer, three incorrect options derived from the human-judgment evaluations of model or human responses to the open-ended riddles (§4.2), and a “cannot determine” option for ambiguous cases. For the 12% instances where fewer incorrect responses were available (i.e., when most responses by humans and models were correct), we used GPT-4 to generate additional distractors, using two in-context examples of visual riddles with incorrect answers. Finally, accuracy is calculated as the proportion of riddles correctly solved by the models, with random guessing yielding a baseline success rate of 20%.
Results
The results in Table 2, indicate that this multiple-choice version is comparably challenging to the open-ended visual riddles, with GPT-4, Gemini-Pro-Vision and Gemini-Pro-1.5 showing the highest accuracies of 45%, 41% and 38%, respectively. Overall, model performance on this task slightly exceeds that on the open-ended task as assessed by human evaluators. Excluding instances where models selected the “cannot determine” option enhances these figures, elevating GPT-4 to 52% and Gemini-Pro-1.5 to 48%. This trend, detailed further in A.7, reveals that models often default to “cannot determine” in the absence of sufficient information. This tendency is mitigated by providing hints and attributions, making GPT-4’s accuracy significantly increase to 69% and 82%, respectively.
Table 2: Three types of multiple choice evaluations: overall accuracy, accuracy excluding instances where the model selected "cannot determine", and accuracy with auxiliary data (hints or attributions).
% Accuracy $\uparrow$ % Cannot % Accuracy w/o + Hint $\uparrow$ + Attribution $\uparrow$
Determine Cannot Determine
Gemini Pro 1.5 38 20 48 66 72
Gemini Pro Vision 41 3 42 62 -
GPT4 45 12 52 69 82
LLaVA-1.6-34B 24 8 26 30 -
LLaVA-1.5-7B 17 0 17 29 -
4.4 Open-ended VQA Automatic Evaluation
The automatic evaluation experiments are designed to assess the capability of vision and language models to accurately judge the correctness of open-ended answers to visual riddles. This evaluation is critical for identifying the most effective approach for automated evaluation, supporting scalable future work on our benchmark, and ensuring integration with the leaderboard intended for widespread use by the research community.
Comparing Auto-Raters to Human Ratings
We evaluate the models in two scenarios: reference-free and reference-based, in the first scenario, models independently assess the correctness of an answer based solely on the image and its associated question, and in the second, models are additionally provided with the ground-truth answer along with the candidate answer, challenging them to validate the provided answer against the correct one. Each auto-rater candidate was evaluated on a subset of two responses provided by other models and humans, excluding its own responses from this evaluation. This subset consists of balanced responses, including one correct and one incorrect. If one of these options was not available, two responses from the same category were selected. We tested the candidate auto-rater models on the visual riddles annotated in §4.2 to determine which model’s performance most closely aligns with human judgments. This alignment is crucial for selecting models suitable for evaluative roles in automated systems.
Table 3: Candidate auto-raters performances compared to human rating.
Judge Accuracy of Judge Prediction
Compared to Human Rating $\uparrow$
Reference-Based Gemini Pro 1.5 87
Gemini Pro Vision 75
GPT4 86
LLaVA -1.6-34b 76
LLaVA -1.5-7b 70
Reference-Free Gemini Pro 1.5 52
Gemini Pro Vision 38
GPT4 51
LLaVA -1.6-34b 43
LLaVA -1.5-7b 35
Results
Table 3 presents the accuracy of the models under two scenarios, with Gemini-Pro-1.5 achieving the highest performance—52% in the reference-free context and 87% in the reference-based context, as measured against human ratings. These results suggest that the reference-free scenario may be less suitable for auto-evaluation, as evidenced by the top score of 52%, indicating a moderate correlation with human judgment. However, the reference-based scenario shows a stronger correlation, with the top two models achieving 87% and 86%. These findings indicate the need for an appropriate judge to assess open-ended answers.
Utilizing the Optimal Auto-Rater to Evaluate Visual Riddles
Having established Gemini-Pro-1.5 as the best auto-rater, we utilize it in a reference-based setting to assess all models on the three open-ended settings detailed in §4.2, i.e., the main task, and its hint- and attribution-assisted variants.
Table 4: Automatic evaluation of open-ended answers by Gemini-Pro-1.5.
Visual Riddles Hints Attribution
% Human % Auto-Rater % Auto-Rater % Auto-Rater
Rating $\uparrow$ Rating $\uparrow$ Rating $\uparrow$ Rating $\uparrow$
Gemini Pro 1.5 40 53 62 29
Gemini Pro Vision 30 34 47 -
GPT4 34 38 61 25
LLaVA -1.6-34b 15 21 16 -
LLaVA -1.5-7b 13 19 30 -
InstructBlip 13 20 28 -
Humans 82 78 - -
Table 4 presents the auto-rating accuracy of various models on the open-ended task, with Gemini-Pro-1.5 as the evaluator. Gemini-Pro-1.5 achieves the highest performance, scoring 53% (40% with human rating) on open-ended questions, 62% with hints, and 29% with attributions. These results suggest that models perform better when provided with hints but not with attributions. Hints improve model performance by directing attention to visual clues, but they do not notably enhance reasoning capabilities. Conversely, when attributions are provided, models struggle to filter through irrelevant details, indicating significant challenges in solving open-ended visual riddles compared to human levels, even with auxiliary information. This highlights ongoing difficulties in improving models’ visual reasoning capabilities and bridging the gap between human and machine understanding in visual interpretation tasks.
5 Analysis
This section explores two key aspects of the Visual Riddles challenge: assessing models’ use of visual cues through comparisons between original and modified images, and examining generative models’ efficacy in replicating images aligned with our riddle prompts for an automatic generation pipeline.
Table 5: Percentage of correctness rates of ground truth answers for original and altered images. The ‘Gap’ column quantifies the reduction in performance due to image modifications, illustrating the challenge of context changes for model accuracy.
Original
[…]
Modified
[…]
Gemini Pro 1.5 74 14 60
Gemini Pro Vision 86 51 35
GPT4 68 15 53
Llava-1.6-34b 93 53 20
Llava-1.5-7b 54 38 16
5.1 Assessing Visual Aspects of Riddles: Ablation Study with Modified Images
To assess models’ preference for specific answers, we selected 72 images from our benchmark and created modified versions, altering visual cues to invalidate ground truth answers. This study explores whether models base their answers solely on text or consider visual clues. We conducted two rounds of testing: one with original images and one with modified versions. In both instances, we presented the model with the question, the original ground truth answer, and asked,“Is the answer correct?” (see Fig. 4). Post-image modification, known ground truth answers became incorrect, make the evaluation to a binary assessment. For each model, we calculated the percentage of correctness rates on each type of image and the gap. A high gap indicates that the model identified the answer as correct for the original image and incorrect for modified one. Table 5 reveals models’ preference for specific answers despite changes in critical image elements and their struggle to identify correct answers, especially with modified images. Gemini-Pro-1.5 excels with a 60% gap, while others below 53%.
Figure 4: Modified images ablation study: a demonstration of the process where the model evaluates an answer’s validity using two scenarios: one with the original image and another with a modified image that alters the visual clue, affecting the correctness of the original ground truth answer.
5.2 The Visual Riddles Prompt Set: A Challenge for Text-to-Image Models
This section shifts our focus from evaluating VLMs and LLMs to showcasing the unique challenge posed by the Visual Riddles prompts to text-to-image models. Designing images for the Visual Riddles dataset, with their specific and often subtle visual clues, proved surprisingly difficult. To quantify this difficulty, we attempted to reproduce 100 visual riddles using five popular open-source diffusion models: SDXL [17], SDXL-Turbo [18], LCM-SDXL [19], SD-1.4, and SD-2.1 [20]. Each model generated 100 images based on these prompts, amounting to 500 images in total. However, only 61 (12%) successfully matched the prompts. The most successful model was SDXL-Turbo, achieving a 15% success rate. A complete breakdown of the models’ performance is available in A.8Fig. 5 showcasing several models struggling to capture the nuances embedded in the prompt. This low overall success rate underscores the unique challenge presented by the Visual Riddles prompts, establishing them as a valuable resource for evaluating and advancing text-to-image generation, especially for tasks that demand high precision and the ability to render subtle visual features.
Figure 5: Reproducing visual riddles: all models fail to capture the nuance in the description.
6 Related Work
Our research is closely linked to commonsense reasoning in multimodal models and the evaluation of factuality in both language-only and multimodal settings.
Commonsense reasoning in multimodal models.
Recent progress in vision-language models including models like BLIP2 [21], Flamingo [22], LLaVA [12], GPT4 [15], and Gemini-Pro [8]. These developments have sparked interest in commonsense reasoning [2, 23], leading to complex visual reasoning challenges within the vision-and-language domain. These include specialized tests for understanding associations and analogies [24, 25], interpreting unusual images in WHOOPS! [26], visual abductive reasoning tasks (e.g., Sherlock; [27]), multi-modal humor understanding tasks [3], and world-instructed image-editing tasks requiring commonsense reasoning (e.g., EditWorld; [28]). A related benchmark to Visual Riddles is VisIT-Bench [5], where the authors took existing images and generated challenging questions about them. However, Visual Riddles distinguishes itself by allowing annotators more creative freedom to devise challenging scenarios and generate corresponding images and questions, rather than restricting them exclusively to what appears in the images. This also ensures the images have not been previously seen during pretraining.
Factuality evaluation.
Evaluating factuality has been a focal point in language-only models, particularly for tasks where outputs rely on grounding texts like summarization or machine translation. This domain utilizes reference-based metrics (e.g., Rouge [29], BLEU [30], Bertscore [31], COMET [32]) that compare outputs to a reference, alongside reference-free metrics (e.g., SummaC [33], $Q^{2}$ [34], TRUE [35], FiC [36]) that evaluate outputs against the input texts. Parallel advancements in vision-language tasks have introduced metrics such as Clipscore [37], AVIBench [38], and Tifa [39], with benchmarks like MileBench [40], SeeTRUE [41], BLINK [42], and Vibe-Eval [43] enhancing the rigor of factuality assessments. Our dataset, Visual Riddles, extends these challenges into the realm of visual commonsense, requiring both deep world knowledge and nuanced commonsense reasoning to interpret complex visual cues. This requirement marks a significant step beyond traditional benchmarks, like Encyclopedic VQA [6], which predominantly tests factual knowledge against curated text sources. By demanding high sensitivity to visual subtleties alongside robust commonsense analysis, Visual Riddles offers a uniquely stringent test of multimodal model capabilities.
7 Conclusions and Limitations
Our analysis of the Visual Riddles Challenge shows that state-of-the-art vision-language models face difficulties in interpreting complex visual scenarios that require commonsense reasoning. With an average success rate of 40%, models significantly lag behind human performance, which stands at 82%. This performance gap underscores the ongoing challenge of bridging human and machine understanding in complex visual interpretation tasks. Visual Riddles, though smaller than other benchmarks, was meticulously annotated to ensure a rich and diverse set of commonsense challenges. While we employed a variety of generative tools to mitigate potential biases, it is crucial to acknowledge that relying on these tools might introduce inherent limitations. Although we have made efforts to exclude visual riddles that could be considered offensive, it is still possible that individuals may find certain riddles offensive, despite our best efforts to prevent this. A promising future direction is to explore model performance across multiple images generated from the same prompt, allowing for a more robust evaluation of visual reasoning capabilities.
Appendix A Appendix
The subsequent sections offer essential insights into reproducibility, encompassing detailed explanations and examples regarding model evaluation, analysis, and data collection.
For access to the complete Visual Riddles dataset, visit https://huggingface.co/datasets/visual-riddles/visual-riddles. Comprehensive information about the dataset fields and loading instructions is provided therein.
A.1 Reproduciblity and Resources
We executed all evaluation code on a single Colab notebook, accessible on our website https://visual-riddles.github.io/. For models using APIs like Gemini and GPT-4, we employed CPU resources. However, for models like Llava and InstructBLIP, which required GPU resources, we utilized a single A100 GPU. We utilized a paid plan for models with APIs, enabling us to generate 400 predictions for each task within a timeframe of less than 30 minutes, equivalent to the runtime of Llava and InstructBLIP models used locally with GPUs. Table 6 lists the models specified in the APIs, and how they were used.
Auxiliary Data: Attributions
For each visual riddle containing an attribution field, we extracted the textual content from the webpage URL using Selenium WebDriver (https://www.selenium.dev/). Consequently, prompts containing attributions may be lengthy.
Table 6: APIs model specification
API Model Specified Name Type of Model Used Tasks Used
GPT4 gpt-4-vision-preview VLM Open-Ended, Multiple-Choice (Distractors + Evaluation), Auto-Eval Judge
Gemini-Pro-1.5 gemini-1.5-pro-latest VLM Open-Ended, Multiple-Choice, Auto-Eval Judge, Auto-Rater
Gemini-Pro-1.5 gemini-1.5-pro-latest LLM Caption $\rightarrow$ LLM on Open-Ended
Gemini-Pro-Vision gemini-pro-vision VLM Open-Ended, Multiple-Choice, Auto-Eval Judge
A.2 Image Generation Guidelines
The Visual Riddles benchmark was created by seven designers experienced in generating images using text-to-image models. The annotators were instructed to create images that integrate information with world knowledge and common sense to answer a given textual question. The answer must be grounded in the data presented in the image, making it impossible to respond accurately without comprehending the image and identifying the embedded clues. To generate high-quality images, our designers are given access to advanced text-to-image models, including Midjourney [^1] https://www.midjourney.com/home, Ideogram [^2] https:///www.ideogram.ai, Canva [^3] https://www.canva.com, DALL-E 3 [10], and Stable-Diffusion [11]. Each image must be sufficiently challenging such that at least one of the evaluation models (e.g., Gemini-Pro-1.5 [7], GPT4-turbo-preview [15] and LLaVA-1.6-34B [13]) fails to correctly answer the question based on the provided image. In addition, the designers provided not only the correct answer to the question but also a hint to the image that should guide where to look in the image when trying to solve the riddle (Example in Fig. 1) and, for riddles requiring world knowledge, include attributions to relevant sources (see bottom example in Fig. 1). After creation, each image-question pair undergoes a peer review by at least three other designers to ensure the hint’s clarity and the riddle’s solvability.
A.3 Categories and Difficulty Level: Breakdown and Results
Figure 6: A heatmap depicting how models (Top) and human annotators (Middle) fail to solve visual riddles in the Visual Riddles dataset, visualizing by the color intensity of each cell. The x-axis represents different categories of riddles, while the y-axis shows the difficulty index levels. Red cells highlight particularly challenging areas, with vertical red bands indicating categories that consistently pose difficulties and horizontal red bands confirming the appropriateness of the difficulty index levels assigned. This visualization underscores the diverse range of commonsense knowledge needed to effectively tackle the visual-riddles. Below, there is an example of images related to some of our categories.
To examine how different types of world knowledge and difficulty levels impact model performance, we analyze their correlations to identify particularly challenging combinations for current models. The distribution of the different categories and difficulty level (with 3 - most difficult and 0 - least difficult) are illustrated in Tables 7 and 8.  Fig. 6 presents a heatmap illustrating model failures and human failures across various categories and difficulty levels within the Visual Riddles challenge. Across 16 categories and 4 difficulty levels, our analysis identifies model weaknesses in categories like “Object Counting”, “Temporal Principles”, and “Physical Principles”. Notably, human performance on “physical principles” appears to be even more susceptible to errors compared to the models on difficulty index level of 3. Naturally, models as well as humans also seem to struggle more with the higher difficulty instances (levels 2 and 3 in  Fig. 6), highlighting the most demanding elements of the benchmark for current vision-and-language models. Yet this highlight the gaps between humans and models performances.
Table 7: Distribution of Categories Across Images
Category % Images
Cultural Principles 19.50
Social Norms 15.75
Safety and Survival 8.25
Food-related 7.50
Object Counting 7.00
Health 6.00
Traffic 5.00
Animal Behavior 5.00
Geographic Knowledge 4.50
Entertainment 4.00
Architectural 3.75
Temporal Principles 3.00
Environmental 3.00
Religion Principles 3.00
Physical Principles 2.50
Biological Principles 2.25
Table 8: Distribution of Categories Across Images
Difficulty Level % Images
0 11.10
1 33.25
2 38.25
3 17.50
A.4 Prompts for Different Models
There are several tasks with different prompts:
  1. 1.
    Multiple-Choice VQA: A specific prompt is used for generating distractors when there are insufficient incorrect answers using GPT-4. Additionally, for multiple-choice VQA, we evaluate the models’ performance across three settings: (1) given only the image, the question, and five possible answers (with only one correct answer), (2) given the image, the question, the possible answers, and a hint, and (3) given the image, the question, the possible answers, and an attribution. The prompts structure is outlined in Table 9.
  2. 2.
    VQA Automatic Evaluation: To find the best judge (auto-rater), we evaluate the models in two scenarios - reference-free and reference-based. The prompts structure is outlined in Table 10.
  3. 3.
    Auto-Rater for Open-Ended VQA and Ablation Study with Modified Images: Using the best Auto-Rater we evaluate all models automatic ratings on the Open-Ended VQA task. We also use the same prompt in our analysis we perform ablation study to explores whether models base their answers solely on text or consider visual clues. The prompts structure is outlined in Table 10.
Table 9: Prompts for different models for Multiple choice VQA
Task Prompt
[…]
Table 10: Prompts for different models for Automatic Evaluation
Task Prompt
[…]
A.5 Open-Ended VQA: Annotators Solve Visual Riddles Guidelines
In order to evaluate how well humans are capable in solving the visual riddles questions, a human response were collected using Amazon Mechanical Turk platform www.mturk.com. To gather human responses for the benchmark’s riddles, we used the Amazon Mechanical Turk platform, paying annotators \$18 per hour. We contacted workers with a proven track record in similar tasks and invited them to a qualification round that began with a review of task guidelines and included solving five riddles of varying difficulty. Following an assessment of their responses and providing personalized feedback, only those demonstrating a strong understanding of the task qualified. Of 14 candidates, 10 proceeded to the actual annotation, where each of the 400 riddles was solved by three annotators. In the guidelines, annotators were presented with five examples. For each example, they were initially shown the visual riddle and subsequently given the solution and the process for solving the question. We recommended that annotators first attempt to solve the riddles on their own before reviewing our provided answers.
Two examples of the guidelines are in  Fig. 7 and  Fig. 8. In the first, answering the question required only common-sense and counting capabilities of objects in the image while in the second example, world knowledge about cultural principles is needed therefore, the workers were expected to search the different items, understand the clues given in the image and only then answer the question.
In  Fig. 9 there is an example of the UI page for annotation of solving a visual riddle.
Figure 7: Example of the guidelines for MTurk annotators for solving a visual riddle as an open ended question. (1) Question: Solving visual riddles composed of images and open-ended questions. (2) Answer: the answer to the question. (3) Explanation & Notes: explanation of why this is the correct answer given the question and the image.
Figure 8: Example of the guidelines for MTurk annotators for solving a visual riddle as an open ended question. (1) Question: Solving visual riddles composed of images and open-ended questions. (2) Answer: the answer to the question. (3) Explanation & Notes: explanation of why this is the correct answer given the question and the image. (4) Attribution: for this image a world knowledge is required therefore - a search in google is helpful in getting the data.
Figure 9: Example of MTurk UI page for answering the question given an image and a question
A.6 Open-Ended VQA: Humans and Models Answers Annotation Evaluation Guidelines
In order to evaluate how well models and humans answered the visual riddle, we used MTurk platform, paying 18\$ per hour. We contacted workers and invited them to a qualification round that began with a review of task guidelines and included solving five riddles of varying difficulty. Following an assessment of their responses and providing personalized feedback, only those with a strong understanding of the task moved on to the actual annotation phase. Of 9 candidates, 6 proceeded to the actual annotation phase.
During annotation, In the LVLM and the Caption $\rightarrow$ LLM cases, each annotator was provided with the image, the question, and the correct answer, along with the candidate answers from six different models and three humans (the human answers were obtained as described in §4.2). In the same way (image, question and ground truth answer), we annotated the candidate answers for human (oracle) and Gemini-Pro-1.5 captions from the best model as described in §4.2.
In the guidelines, annotators were presented with five examples. For each example, they were initially shown the image, the question, the gold-answer, and the candidate answers. Afterwards, they could see which answers were correct and the reasoning behind the correctness or incorrectness of each candidate answer. Marking a candidate answer as correct required that it not only accurately answered the question based on the given answer but also contained no hallucinations.
An example of the guidelines as well as a UI page is in Fig 10. Using these annotations as mentioned in §4.3, we compose the Multiple-Choice VQA prompts by sampling three answers annotated as incorrect. If there are fewer than three, we generate additional incorrect answers. We then include the ground-truth answer provided by the riddle designer, along with the "cannot determine" option.
Figure 10: Example of guidelines for MTurk annotators evaluating different answers to the visual riddle. (1) Annotator UI page displaying the visual riddle, the correct answer, and the candidate answers. (2) Solution: Identification of which candidate answers correctly solve the visual riddle. (3) Explanation & Notes: Detailed reasoning for the correctness or incorrectness of each candidate answer (only explanations for three candidates are shown due to space constraints).
A.7 Multiple Choice VQA Analysis: Model Hesitation
As mentioned in §4.3, a notable occurrence was the frequent selection of the "cannot determine" option by some models. Further analysis excluding these instances revealed improvements in comparison to the results with this option (see Table 2), with GPT-4 achieving 52% accuracy instead of 45%, and Gemini-Pro-1.5 reaching 48% accuracy instead of 38%, respectively. These results suggest that some models hesitate to answer certain questions, opting for the "cannot determine" option when lacking sufficient information to decide. Additionally, as shown in Table 11, providing models with auxiliary information via hints and attributions reduces their likelihood of selecting the "cannot determine" option. For example, Gemini-Pro-1.5 achieves an accuracy rate of 74% with a hint, while GPT-4 achieves an accuracy rate of 84% with an attribution.
Table 11: Percentage of “Cannot Determine” Answers
% Cannot + Hint + Attribution
Determine % Cannot Determine % Cannot Determine
Gemini Pro 1.5 20 12 9
Gemini-Pro-Vision 3 2
GPT4 12 3 3
LLaVA-1.6-34B 8 10
LLaVA-1.5-7B 0 0
Table 12: Accuracies Excluding “Cannot Determine” Answer
% Accuracy w/o + Hint % Accuracy + Attribution % Accuracy
Cannot Determine w/o Cannot Determine w/o Cannot Determine
Gemini Pro 1.5 48 74 79
Gemini-Pro-Vision 42 64
GPT4 52 71 84
LLaVA-1.6-34B 26 34
LLaVA-1.5-7B 17 29
A.8 The Visual Riddles Prompt Set: A Challenge for Text-to-Image Models
n this section, we present the complete breakdown of the models’ performance in generating images that fit the visual-riddles prompts. Table 13 presents the full results, indicating that models struggle to generate images that include delicate hints hidden within them, with the best-generating model, SDXL-Turbo [18], creating only 15% of images that match the given prompt.
Table 13: Model success rates
% Generation
Success $\uparrow$
SD-1.4 7
SD-2.1 12
SDXL-LCM 12
SDXL 14
SDXL-Turbo 15
Appendix B Discussion any potential negative societal impacts Visual Riddles
While Visual Riddles offers a unique platform for enhancing visual reasoning and commonsense understanding, it also presents potential challenges that merit careful consideration. One concern is the inadvertent reinforcement of biases. Despite rigorous efforts to design inclusive and culturally neutral visual riddles, the possibility remains that some content might unintentionally reflect or amplify societal stereotypes. Moreover, the complexity of certain riddles could disadvantage users with specific cognitive or sensory impairments, thereby limiting their participation and representation in the research facilitated by this benchmark.
Another limitation is the reliance on automated evaluation methods. Such methods may not fully capture the depth of human reasoning and interpretation, potentially affecting the robustness and transparency of the evaluations. This might inadvertently prioritize certain types of reasoning over others, skewing the development and assessment of AI systems.
To address these issues, we are committed to a continuous review process involving diverse stakeholders to help identify and mitigate any biases or exclusions. This process will include refining the riddles and improving the evaluation methods to ensure they are as inclusive and representative as possible. Additionally, ongoing adjustments will be made to the evaluation protocols to enhance their ability to assess nuanced and complex responses, thereby ensuring that the benchmark remains a fair and effective tool for advancing AI research in an ethically responsible manner.
Appendix C Designer Consent
We acknowledge and extend our gratitude to all designers who contributed to the benchmark. The credit for each visual riddle is included as part of our dataset. All designers have consented to contribute their creations to this research.